OpenStudy (anonymous):

When looking at a rational function, Charles and Bobby have two different thoughts. Charles says that the function is defined at x = −2, x = 3, and x = 5. Bobby says that the function is undefined at those x values. Describe a situation where Charles is correct, and describe a situation where Bobby is correct. Is it possible for a situation to exist where they are both correct? Justify your reasoning.

OpenStudy (anonymous):

so you up on rational functions?

OpenStudy (anonymous):

YES I CAME UP WITH THE EQUATION Y=X

OpenStudy (anonymous):

rational functions are generally the ratio of poynomials, like \[f(x)=\frac{ p(x) }{ q(x) }\]where f(x) is the rational function, p(x) and q(x) are both polynomials. so long as the polynomials have only non-negative, integer exponents, the rational polynomial will be defined for all values of x such that the denominator (q(x)) is NOT 0. f(x) is NOT defined when q(x) = 0. for example:\[f(x) = \frac{ 1 }{ x-2 }\] what makes the denominator 0?

OpenStudy (anonymous):

i dont exactly know, can you walk me through the problem please

OpenStudy (anonymous):

what value would x have to be in order to make the denominator 0?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

yep, awesome! so because dividing by 0 isn't allowed, x = 2 is not defined for that function

OpenStudy (anonymous):

ok can you help find a function that works for both their statements

OpenStudy (anonymous):

there isn't a way a function can be both defined and undefined at the same point. it's either one or the other.

OpenStudy (anonymous):

so how to i justify that cant be both i have to have an explanation

OpenStudy (anonymous):

can a gum be both sugar-free and have sugar in it?

OpenStudy (anonymous):

i mathmatical reason like an example of a function that wont work for both

OpenStudy (anonymous):

make one up for each. for one that isn't defined at x = -2, move the -2 to the other side => x+2=0 then put it on the bottom => 1/(x+2) and now you have a rational function that isn't defined when x=-2. add the other ones using multiplcation.... \[\frac{ 1 }{ (x+2)(x-3) }\]and now you have a function that isn't defined at x=-2 and isn't defined at x=3. get the idea? for a function that is defined at those point, simply put them in the numerator... (x+2)(x-3) is a function defined at x=-2 and is defined at x = 3.

OpenStudy (anonymous):

it has to be the same equation for both

OpenStudy (anonymous):

let me reiterate, there is NO function that is both defined and undefined at the same point.

OpenStudy (anonymous):

ino but give me a randon equation so i can show how they cant both be right

OpenStudy (anonymous):

hello??